OpenStudy (anonymous):

y=(x-2)/(x-5) solve for x

OpenStudy (anonymous):

\[y=\frac{x-2}{x-5}\] \[y(x-5)=x-2\] \[xy-5y =x-2\] \[xy-x=5y-2\] \[x(y-1)=5y-2\] \[x=\frac{5y-2}{x-1}\]

OpenStudy (anonymous):

y-1

OpenStudy (anonymous):

@satellite denominator malfunction.

OpenStudy (anonymous):

@satellite73